### v2

# Local limits of random spanning trees in random environment

Abstract.
We study the edge overlap and local limit of the random spanning tree in random environment (RSTRE) on the complete graph with $n$ vertices and weights given by $\exp(-\beta\omega_{e})$ for $\omega_{e}$ uniformly distributed on $[0,1]$. We show that for $\beta$ growing with $\beta=o(n/\log n)$, the edge overlap is $(1+o(1))\beta$, while for $\beta$ much larger than $n\log^{2}n$, the edge overlap is $(1-o(1))n$. Furthermore, there is a transition of the local limit around $\beta=n$. When $\beta=o(n/\log n)$ the RSTRE locally converges to the same limit as the uniform spanning tree, whereas for $\beta$ larger than $n\log^{\lambda}n$, where $\lambda=\lambda(n)\rightarrow\infty$ arbitrarily slowly, the local limit of the RSTRE is the same as that of the minimum spanning tree.
Key words and phrases:
disordered system, uniform spanning tree, minimum spanning tree, random graphs
2020 Mathematics Subject Classification:
Primary: 60K35; Secondary: 82B41, 82B44, 05C05
1. Introduction
Let $G=(V,E)$ be the complete graph on the vertex set $V={1,\ldots,n}$, and let $\omega=(\omega_{e})_{e\in E}$ be i.i.d. random variables uniformly distributed on $[0,1]$. The random spanning tree in random environment (RSTRE) on $G$ is defined as follows. Given $(\omega_{e})_{e\in E}$ and $\beta\geq 0$, we let $\textup{{w}}(e)=\exp(-\beta\omega_{e})$ and define the (weighted) uniform spanning tree measure on $\mathbb{T}(G)$, the set of spanning trees on $G$, as
$\boldsymbol{\mathrm{P}}_{n,\beta}^{\omega}(\mathcal{T}=T):=\frac{1}{Z_{\beta}^{\omega}}\prod_{e\in T}\textup{{w}}(e)=\frac{1}{Z_{\beta}^{\omega}}\prod_{e\in T}\exp(-\beta\omega_{e}),$ (1.1)
where $Z_{\beta}^{\omega}$ is the normalization constant, also known as the partition function. We write $\boldsymbol{\mathrm{E}}^{\omega}_{n,\beta}$ for the expectation w.r.t. $\boldsymbol{\mathrm{P}}_{n,\beta}^{\omega}$, and we will often drop the sub- and superscripts when there is no ambiguity. Instead of studying the quenched law (1.1) for fixed realizations of $\omega$, we will study the averaged (over $\omega$) law
^$\widehat{\mathbb{P}}_{n,\beta}(\cdot)=\widehat{\mathbb{P}}(\cdot)=\mathbb{E}\big{[}\boldsymbol{\mathrm{P}}(\cdot)\big{]}$
with corresponding expectation $\widehat{\mathbb{E}}$.
In [9], the authors showed that for arbitrary distributions, but fixed $\beta$, the diameter of typical trees on sequences of either bounded degree expanders or boxes in $\mathbb{Z}^{d}$, $d\geq 5$, grows asymptotically as $|V|^{1/2+o(1)}$. In [10] the same authors further analyzed the diameter of a typical tree on the complete graph with the environment being distributed as i.i.d. uniforms on $[0,1]$. They show the existence of two regimes: (1) a low disorder regime $\beta\leq n^{1-o(1)}$, where the diameter is of the same order as the diameter of the uniform spanning tree (UST); (2) a high disorder regime $\beta\geq n^{4/3+o(1)}$, where the diameter is comparable to the diameter of the random minimum spanning tree (MST). In this paper, we shall study local observables for the RSTRE with the same setting as in [10].
Let ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ be two spanning trees sampled independently according to $\boldsymbol{\mathrm{P}}^{\omega}_{n,\beta}$. The first observable that we shall study is the expected edge overlap
$\mathcal{O}(\beta):=\boldsymbol{\mathrm{E}}^{\omega}_{n,\beta}\boldsymbol{\mathrm{E}}^{\omega}_{n,\beta}\big{[}|{\mathcal{T}}\cap{\mathcal{T}}^{\prime}|\big{]}=\sum_{T,T^{\prime}\in\mathbb{T}(G)}|T\cap T^{\prime}|\,\boldsymbol{\mathrm{P}}^{\omega}_{n,\beta}({\mathcal{T}}=T)\boldsymbol{\mathrm{P}}^{\omega}_{n,\beta}({\mathcal{T}}=T^{\prime}),$ (1.2)
where we identify the trees with their edges and let $\beta=\beta(n)$ depend on $n$. Our first main result is the following, where we say that a sequence of events $A_{n}$ holds with high probability if ^$\widehat{\mathbb{P}}(A_{n})\rightarrow 1$ as $n$ tends to infinity.
Theorem 1.1.
Let ${\mathcal{T}}$ be the RSTRE on the complete graph with $n$ vertices and random environment distributed as uniform random variables on $[0,1]$. If $0<\beta=\beta(n)\ll n/\log n$, then with high probability
$\mathcal{O}(\beta)=(1+o(1))\beta\frac{1-e^{-2\beta}}{(1-e^{-\beta})^{2}}.$ (1.3)
If on the other hand $\beta\gg n(\log n)^{2}$, then with high probability
$\mathcal{O}(\beta)=(1-o(1))n.$ (1.4)
Remark 1.2.
As the probability of an edge being in the (unweighted) UST on the complete graph (which corresponds to $\beta=0$ in (1.1)) is $2/n$ and is independent of $\omega$, the expected overlap of the UST is $2(n-1)/n$. On the other hand, for the MST the overlap is trivially $(n-1)$.
Our second result is the local (weak) convergence of the RSTRE to either the UST local limit or the MST local limit depending on the strength of the disorder $\beta=\beta(n)$. We refer to Section 2.4, and in particular to (2.7), for a characterization of local (weak) convergence of the RSTRE.
Theorem 1.3.
Let ${\mathcal{T}}$ be the RSTRE on the complete graph with $n$ vertices and random environment distributed as uniform random variables on $[0,1]$. If $\beta=\beta(n)\ll n/\log$, then under the averaged law $\widehat{\mathbb{P}}$ we have the local convergence
[…]
where ${\mathcal{T}}(1)$ is ${\mathcal{T}}$ rooted at $1$ and $\mathcal{P}$ is a Poisson(1) branching process conditioned to survive forever. If on the other hand $\beta=\beta(n)\geq n(\log n)^{\lambda}$, where $\lambda=\lambda(n)\rightarrow\infty$ arbitrarily slowly, then there exists a (random) rooted tree $(\mathcal{M},o)$ such that under the averaged law the local convergence
[…]
holds.
To construct $\mathcal{P}$, the Poisson(1) branching process conditioned to survive forever, we first consider an infinite backbone $x_{0},x_{1},x_{2},\ldots$ started at the root $x_{0}=o$, and then attach independent Poisson(1) branching processes to each vertex $x_{i}$. For a large class of graphs, this is the local limit of the UST (see e.g. [4] and [11]). The tree $\mathcal{M}$ is the local limit of the MST on the complete graph as obtained in [1]. It may be constructed by starting an invasion percolation process at every vertex of the Poisson weighted infinite tree rooted at $o$, and then taking the union of all clusters containing $o$. We refer to [1] for more details.
Theorem 1.3 shows that there is a sharp cut-off for the local limit when $\beta=n^{\gamma}$ and $\gamma$ crosses the critical value $\gamma_{c}=1$. This stands in contrast to [10, Conjecture 1.3], where the diameter is conjectured to smoothly interpolate between $n^{1/2}$ and $n^{1/3}$ when $\gamma\in[1,4/3]$. It would be interesting to study what the local limit (if it exists) should be when $\beta$ is of order $n$. Both the proofs of Theorem 1.1 and Theorem 1.3 will make use of the fact that we are able to give sharp bounds on the effective resistance (see Definition 2.1).
2. Preliminaries
In this section, we first recall the relevant notation for electric networks, following mainly the exposition of Chapter 2 of [8]. Then we discuss a coupling of the random environment and of the MST to an Erdős-Rényi random graph, and lastly, we recall the notion of local convergence of graphs needed for Theorem 1.3.
2.1. Electric Networks and UST
Let $(G,\textup{{w}})$ be a finite graph with positive weights $\textup{{w}}(\cdot)$ on the undirected edges $e=(u,v)=(v,u)$. For each edge $e\in E$ fix an arbitrary orientation and let $e^{-}$ and $e^{+}$ be the starting and endpoints of $e$. It is useful to think of $E$ containing each edge in both orientations. Given disjoint vertex sets $A,B\subset V$, a unit flow $\vartheta:E\rightarrow\mathbb{R}$ from $A$ to $B$ is an antisymmetric function that has net flow out of $A$ equal to $1$, net flow out of $B$ equal to $-1$, and zero net flow out of any vertices not in $A\cup B$. That is, $\vartheta$ satisfies
  1. (1)
    $\vartheta(e^{-},e^{+})=-\,\vartheta(e^{+},e^{-})$;
  2. (2)
    $\sum_{e^{-}=x}\vartheta(e)=0$ for all $x\not\in A\cup B$;
  3. (3)
    and
    $\sum_{a\in A}\sum_{e^{-}=a}\vartheta(e)=1=-\sum_{b\in B}\sum_{e^{-}=b}\vartheta(e).$
The energy of a flow $\|\vartheta\|^{2}_{r}$ is defined as
$\|\vartheta\|^{2}_{r}:=\frac{1}{2}\sum_{e\in E}\frac{1}{\textup{{w}}(e)}\vartheta(e)^{2},$ (2.1)
where each edge appears with both orientations in the sum.
Definition 2.1 (Effective resistance).
Let $A,B\subset V$ be two disjoint sets of vertices. We define the effective resistance as
${\rm R}_{\rm eff}^{(G,\textup{{w}})}(A\leftrightarrow B):=\min\left\{\|\vartheta\|^{2}_{r}\,:\,\vartheta\text{ is a unit flow from }A\text{ to }B\right\},$
and will often drop the superscripts when there is no ambiguity. The minimizing flow is called the unit current flow.
One useful connection between electric networks and USTs is Kirchhoff’s formula.
Theorem 2.2 (Kirchhoff’s formula [8, Section 4.2]).
Given a weighted graph $(G,\textup{{w}})$, the probability that an edge $(u,v)$ is in the (weighted) uniform spanning tree is
$\boldsymbol{\mathrm{P}}\big{(}(u,v)\in\mathcal{T}\big{)}=\textup{{w}}(u,v){\rm R}_{\rm eff}(u\leftrightarrow v).$ (2.2)
Applying Kirchhoff’s formula to the overlap $\mathcal{O}(\beta)$ from (1.2) shows that
$\mathcal{O}(\beta)$ $=\boldsymbol{\mathrm{E}}^{\otimes 2}\Big{[}\sum_{(u,v)\in E}1_{(u,v)\in{\mathcal{T}},(u,v)\in{\mathcal{T}}^{\prime}}\Big{]}$
$=\sum_{(u,v)\in E}\boldsymbol{\mathrm{P}}\big{(}(u,v)\in{\mathcal{T}}\big{)}^{2}=\sum_{(u,v)\in E}\textup{{w}}(u,v)^{2}{\rm R}_{\rm eff}(u\leftrightarrow v)^{2}.$ (2.3)
In fact, the theory of electric networks allows one to calculate the probability of any collection of edges being in the tree. We formalize this in the following definition and theorem.
Definition 2.3 (Transfer-impedance matrix).
Recall that we fixed an (arbitrary) orientation of the edges. Let $v_{e}(\cdot)$ be the voltages on the vertices when a unit current flow $\vartheta_{e}(\cdot)$ is sent from $e^{-}$ to $e^{+}$, i.e. $v_{e}(\cdot)$ is the (weighted) harmonic function on $V\setminus\{e^{-},e^{+}\}$ with boundary values $v_{e}(e^{-})={\rm R}_{\rm eff}(e^{-}\leftrightarrow e^{+})$ and $v_{e}(e^{+})=0$. The transfer-impedance matrix $Y$ is a $|E|\times|E|$ matrix with entries given by
$Y(e,f):=\textup{{w}}(f)\big{[}v_{e}(f^{-})-v_{e}(f^{+})\big{]}=\vartheta_{e}(f^{-},f^{+}).$
The following theorem first appeared in [4].
Theorem 2.4 (Transfer-impedance theorem).
For distinct edges $e_{1},\ldots,e_{k}$
$\boldsymbol{\mathrm{P}}(e_{1},\ldots,e_{k}\in{\mathcal{T}})=det\big{(}Y(e_{i},e_{j})_{1\leq i,j\leq k}\big{)},$ (2.4)
where $Y(e_{i},e_{j})_{1\leq i,j\leq k}$ are entries of the transfer-impedance matrix $Y$.
A direct consequence of the transfer-impedance theorem is that for $e\neq f$
$Y(e,f)=\pm\sqrt{\frac{\textup{{w}}(f)}{\textup{{w}}(e)}}\sqrt{\boldsymbol{\mathrm{P}}(e\in{\mathcal{T}})\boldsymbol{\mathrm{P}}(f\in{\mathcal{T}})-\boldsymbol{\mathrm{P}}(e,f\in{\mathcal{T}})}.$ (2.5)
This means that the UST measure is determined, up to signs, by its one- and two-dimensional marginal distribution on the edges.
2.2. Coupling to a random graph
For $p\in[0,1]$ let $G_{n,p}$ be the Erdős-Rényi random graph obtained by independently keeping each edge of the complete graph with probability $p$ and deleting it otherwise. We shall construct $G_{n,p}$ by coupling the random graph to the i.i.d. uniforms $(\omega_{e})_{e\in E}$ by letting
$e\in G_{n,p}\iff\omega_{e}\leq p.$
Further, denote $\mathcal{C}_{1}(p),\mathcal{C}_{2}(p),\ldots$ the largest connected components ordered in decreasing size. The following theorem gives an upper bound on the size of the second-largest component for a wide range of $p$.
Theorem 2.5.
There exists constants $L,\eta>0$ such that for $\epsilon\geq n^{-1/4}$ with $\epsilon=o(1)$ and
$p\not\in\Big{[}\frac{1-\epsilon}{n},\frac{1+\epsilon}{n}\Big{]},$
we have
$\mathbb{P}\Big{(}\big{|}\mathcal{C}_{2}(p)\big{|}\geq L\log(n)\epsilon^{-2}\Big{)}=O(n^{-\eta}).$ (2.6)
Proof Sketch.
The case $0\leq pn\leq 1-\epsilon$ follows by comparison to a (slightly) subcritical branching process (see e.g. Theorem 4.4 in [12] and the proof thereof). For the regime $pn\geq 1+\epsilon$, one first shows that there are no components of intermediate size (see e.g. Corollary 4.13 of [12]), and then one shows that the giant component is unique (see e.g. Lemma 4.14 of [12] or Lemma 3 in [7]). ∎
For $pn=\log n+c$ the probability of a random graph being connected starts to converge to $1$ as $c$ increases. We will use the following theorem, and refer to Theorem 9 in Chapter 7 of [3] for a proof.
Theorem 2.6.
Let $pn=\log n+\kappa(n)$ for $\kappa(n)\rightarrow\infty$. Then for $n$ large enough
$\mathbb{P}(G_{n,p}\textrm{ is not connected}\,)\leq 4e^{-\kappa(n)}$
2.3. Minimum spanning tree
Given distinct weights $\textup{{w}}(\cdot)$ on the edges, the minimum spanning tree (MST) is the unique tree that minimizes
$\sum_{e\in T}\textup{{w}}(e)$
over all $T\in\mathbb{T}(G)$, the set of spanning trees in $G$, and we write $M=M(\textup{{w}})$ for this tree. By letting $\textup{{w}}(e)=\omega_{e}$, we can couple the MST to the environment so that the tree with the largest probability in (1.1) is exactly the MST. This in particular also gives a coupling of the MST to the random graph process defined in the previous section. By using Kruskal’s algorithm (see e.g. the introduction of [1]) to construct the MST, one can see that under this coupling the vertex sets of the connected components of $G_{n,p}$ and $G_{n,p}\cap M$ coincide. This allows us to prove the following lemma as a consequence of Theorem 2.6.
Lemma 2.7.
Let $M$ be the random MST of the complete graph with $\textup{{w}}(e)=\omega_{e}$ where $(\omega_{e})_{e\in E}$ are i.i.d. uniforms on $[0,1]$. Denote by $\deg_{M}(v)$ the degree of $v$ in $M$, then
$\mathbb{P}\big{(}\exists v\in V:\deg_{M}(v)>60\log n\big{)}=O(n^{-4}).$
Proof.
Let $p=5\log n/n$ and $d=60\log n$. If $G_{n,p}$ is connected, then no other edges will be added to the MST in the coupling above, and the maximum degree of a vertex in $M$ is bounded by the maximum degree of a vertex in $G_{n,p}$. Hence, by Theorem 2.6
$\mathbb{P}\big{(}\exists v\in V:\deg_{M}(v)>d\big{)}$ $\leq\mathbb{P}\big{(}\exists v\in V:\deg_{G_{n,p}}(v)>d\big{)}+\mathbb{P}\big{(}G_{n,p}\textrm{ is not connected}\,\big{)}$
$\leq n\binom{n-1}{d}p^{d}+O(n^{-4})\leq n\cdot\Big{(}\frac{e\cdot 5\log n}{d}\Big{)}^{d}+O(n^{-4})$
$\leq\exp\Big{(}d\big{(}2+\log 5+\log\log n-\log d\big{)}\Big{)}+O(n^{-4})$
$\leq\exp\big{(}d(4-\log 60)\big{)}+O(n^{-4})=O(n^{-4}).\qed$
2.4. Local Limits
For $v\in V(G)$ the pair $(G,v)$ is called a rooted graph with root $v$. We say that two rooted graphs $(G,v)$ and $(G^{\prime},v^{\prime})$ are isomorphic if there exists a graph isomorphism $\phi:V(G)\rightarrow V(G^{\prime})$ between $G$ and $G^{\prime}$ fixing the roots, i.e. there exists a bijection $\phi$ between the vertex sets such that $\phi(v)=v^{\prime}$ and $(x,y)\in E(G)$ if and only if $(\phi(x),\phi(y))\in E(G^{\prime})$, and in this case we write $(G,v)\simeq(G^{\prime},v^{\prime})$. For $r=1,2,\ldots$ denote by $B_{G}(v,r)$ the closed ball (viewed as a rooted subgraph) of radius $r$ centered at $v$ in $G$. Let $(G,v)$ be a (random) rooted graph with law $\mu$, and consider a random rooted graph sequence $(G_{n},v_{n})_{n\in\mathbb{N}}$ with corresponding laws $(\mu_{n})_{n\in\mathbb{N}}$, where $(v_{n})_{n\in\mathbb{N}}$ are vertices drawn uniformly from the vertex sets of $(G_{n})_{n\in\mathbb{N}}$. Local (weak) convergence of $G_{n}$ to $G$ may be characterized by the condition that for all $r\geq 1$ and every finite rooted graph $(H,o)$
$\lim\limits_{n\rightarrow\infty}\mu_{n}\Big{(}B_{G_{n}}(v_{n},r)\simeq(H,o)\Big{)}=\mu\Big{(}B_{G}(v,r)\simeq(H,o)\Big{)}.$
We refer to [2] for a more formal introduction.
As the complete graph is transitive and the edge weights are i.i.d., the edge weights form an exchangeable sequence of random variables. In particular, the averaged law is invariant under any permutation of the vertices so we may as well fix the random vertex $v_{n}=1$. In particular, the local weak convergence of the RSTRE to some rooted graph $(\mathcal{S},v)$ with law $\mu$ is equivalent to
^$\widehat{\mathbb{P}}_{n,\beta}\Big{(}B_{{\mathcal{T}}}(1,r)\simeq(H,o)\Big{)}\xrightarrow{n\rightarrow\infty}\mu\Big{(}B_{\mathcal{S}}(v,r)\simeq(H,o)\Big{)},$ (2.7)
[…]
Recall from Theorem 1.3 that $(\mathcal{P},o)$ is the Poisson(1) branching process conditioned to survive forever with root $o$. The following result was first proven in [5].
Theorem 2.8.
The UST on the complete graph converges locally to $(\mathcal{P},o)$.
As mentioned in the introduction, since the work of [5], the local convergence has been extended to other families of graphs in e.g. [4] and [11]. For instance, the authors of [11] show that the above theorem holds for any (almost) $d$-regular graph sequence with degrees diverging as the number of vertices in the graph increases.
The local limit of the MST has been studied less intensively. Nonetheless, in [1] it was proved that the MST on the complete graph converges to a local limit, and by showing that the limit has cubic volume growth compared to the quadratic volume growth of $\mathcal{P}$, that the limiting object is different than $\mathcal{P}$. We state their result in the following theorem, and note that in [1] the convergence is stronger by also taking into account the edge weights of the tree.
Theorem 2.9.
There exists a random rooted tree $(\mathcal{M},o)$ such that the MST converges locally to $(\mathcal{M},o)$.
3. Low Disorder
In this section, we prove Theorem 1.1 and Theorem 1.3 for $\beta=\beta(n)\ll n/\log n$. We first prove lower bounds on the effective resistance using concentration arguments in Section 3.1. Then we prove matching upper bounds in Section 3.2 by constructing a good flow on the weighted graph. The concentration of the effective resistance will then allow us to conclude Theorem 1.1 and, using the method of tree moments developed in [4], we may also prove Theorem 1.3.
3.1. Lower Bounds on the Effective Resistance
We briefly recall a concentration inequality of [10] that will be useful for the lower bounds of the effective resistance. We consider the random weighted graph $(G,\textup{{w}})$ with weights given by
$\textup{{w}}(e)=\exp(-\beta\omega_{e})$
where $(\omega_{e})_{e\in E}$ are i.i.d. uniforms on $[0,1]$, and the mean is given by
$\mu=\mu(\beta)=\frac{1-e^{-\beta}}{\beta}.$
Applying the Berstein inequality as in [10, Section 3.1], we obtain that for $S_{m}$ a sum of $m$ edge weights, $\beta\geq 1$ and any $0<\delta\leq 1$ we have
$\mathbb{P}\big{(}|S_{m}-m\mu|\geq\delta m\mu\big{)}\leq 2\exp\Big{(}-\frac{\delta^{2}m}{9\beta}\Big{)}.$ (3.1)
For simplicity, we shall assume from now on that $\beta\geq 1$, and remark that the concentration arguments follow similarly for $0\leq\beta\leq 1$.
For two edges $e,f$ let $|e\cap f|$ be the number of endpoints that they share. The following lemma will follow quite easily from the concentration inequality (3.1) and the electric network tools developed in Section 2.1. We shall denote by $G/f$ the graph obtained by contracting the endpoints of $f$ into a single vertex.
Lemma 3.1.
Assume $\beta=\beta(n)\ll n/\log n$ and fix $K>0$. Then the following holds with probability at least $1-O(n^{-K})$:
  1. (i)
    for all $u,v\in V$
    ${\rm R}_{\rm eff}(u\leftrightarrow v)\geq\frac{2(1-o(1))}{n\mu};$ (3.2)
  2. (ii)
    for all edges $e,f$ with $|e\cap f|=0$
    ${\rm R}_{\rm eff}^{G/f}(u\leftrightarrow v)\geq\frac{2(1-o(1))}{n\mu};$ (3.3)
  3. (iii)
    and for all edges $e=(u,v)$ and $f$ with $|e\cap f|=1$
    ${\rm R}_{\rm eff}^{G/f}(u\leftrightarrow v)\geq\frac{3(1-o(1))}{2n\mu}.$ (3.4)
Proof.
From Rayleigh’s monotonicity principle (see [8, Section 2.4]), it follows that the effective resistance on $G$ is at least the effective resistance on the graph $G^{\prime}$ where every edge between two vertices $x,y\not\in\{u,v\}$ is contracted (in particular if $|e\cap f|=0$ the edge $f$ is also contracted). On $G^{\prime}$ the series and parallel law (see [8, Section 2.3]) give that
$\frac{1}{\textup{{w}}(u,v)+\Big{(}\big{(}\sum_{y\neq u,v}\textup{{w}}(u,y)\big{)}^{-1}+\big{(}\sum_{y\neq u,v}\textup{{w}}(y,v)\big{)}^{-1}\Big{)}^{-1}}={\rm R}_{\rm eff}^{G^{\prime}}(u\leftrightarrow v)\leq{\rm R}_{\rm eff}(u\leftrightarrow v).$ (3.5)
Fix $M>1$, then by (3.1) with $\delta=3\sqrt{M\beta\log n/n}=o(1)$, and using that $\beta\ll n/\log n$, the inequality
$\sum_{y\neq u,v}\textup{{w}}(u,y)\leq(1+o(1))n\mu$
holds with probability at least $1-O(n^{-M})$. Furthermore, as trivially $\textup{{w}}(i,j)\leq 1\ll n/\beta$, we may easily deduce inequality (3.2) and (3.3).
For inequality (3.4) assume w.l.o.g. that $f=(u,x)$ with $x\neq v$. Repeating the same contraction argument, and additionally contracting $f=(u,x)$, gives again by the series and parallel law that
$\frac{1}{\textup{{w}}(u,v)+\textup{{w}}(x,v)+\Big{(}\big{(}\sum_{y\neq u,v,x}\textup{{w}}(u,y)+\textup{{w}}(x,y)\big{)}^{-1}+\big{(}\sum_{y\neq u,v}\textup{{w}}(y,v)\big{)}^{-1}\Big{)}^{-1}}\\ ={\rm R}_{\rm eff}^{G^{\prime}}(u\leftrightarrow v)\leq{\rm R}_{\rm eff}^{G/f}(u\leftrightarrow v).$
Similar concentration arguments as for (3.5), show that (3.4) holds with probability at least $1-O(n^{-M})$. Choosing $M$ large enough and applying a union bound finishes the proof. ∎
3.2. Upper Bounds on the Effective Resistance
To prove the analogous upper bounds of Lemma 3.1, we will first show upper bounds on the effective resistance on $G_{n,p}$, and then extend these to the weighted graph $(G,\textup{{w}})$ via the coupling in Section 2.2.
Lemma 3.2.
Suppose $\log n\ll pn\ll\sqrt{n}$ and let $\epsilon=\epsilon(n)>0$ be such that $\sqrt{\log n/pn}\ll\epsilon<1/2$. There exists some universal constant $C>0$ such that for any $K>0$ the following holds with probability at least $1-O(n^{-K})$:
  1. (i)
    for all $u,v\in V$
    ${\rm R}_{\rm eff}^{G_{n,p}}(u\leftrightarrow v)\leq\frac{2(1+C\epsilon)}{pn};$ (3.6)
  2. (ii)
    and for all edges $e=(u,v)$, $f=(u,x)$ with $|e\cap f|=1$
    ${\rm R}_{\rm eff}^{G_{n,p}/f}(u\leftrightarrow v)\leq\frac{3(1+C\epsilon)}{2pn}.$ (3.7)
We remark that inequality (3.6) is essentially Proposition 3.1 in [6] with the additional constraint that $\sqrt{\log n/pn}\ll\epsilon$, ensuring that with high probability the degrees in the random graph are of the correct size, and hence we will not prove it. The proof in [6] uses a branching type process to show that there are many different paths between the neighbors of $u$ and the neighbors $v$, which give rise to a flow that captures the behavior of the effective resistance. It is possible to adapt the proof to also show (3.7) by noting that the degree of the first vertex is essentially twice as large as in the non-contracted case. However, instead of repeating many parts of the proof, we will make use of the following lemma.
Lemma 3.3 (Lemma 3.5 in [11]).
Let $G$ be a weighted graph and let $v_{1},\ldots,v_{k}$ be distinct vertices with $k\geq 3$. Let $x>0$ be given and $\epsilon>0$ satisfy $\epsilon\leq\frac{x}{32k}$. Assume that for any distinct $v_{i},v_{j}$ one has
$\big{|}{\rm R}_{\rm eff}(v_{i}\leftrightarrow v_{j})-x\big{|}\leq\delta.$
Then
$\Big{|}{\rm R}_{\rm eff}(v_{1}\leftrightarrow\{v_{2},\ldots,v_{k}\})-\frac{kx}{2(k-1)}\Big{|}\leq 72k\delta.$
Proof of (3.7) in Lemma 3.2.
Restrict to the event that the bound in (3.6) holds for all pairs of vertices, and that the degree of each vertex is between $(1+\epsilon)pn$ and $(1-\epsilon)pn$. As both these events occur with a probability that we can control well, using the trivial lower bound on the effective resistance
${\rm R}_{\rm eff}(u\leftrightarrow v)\geq\frac{1}{\deg(u)+1}+\frac{1}{\deg(v)+1},$
it follows that for all distinct $u$ and $v$
$\big{|}{\rm R}_{\rm eff}(u\leftrightarrow v)-\frac{2}{pn}\big{|}=O\Big{(}\frac{\epsilon}{pn}\Big{)}$ (3.8)
with probability at least $1-O(n^{-K})$ for all pair of vertices and any fixed $K>0$.
Denote by $\tilde{u}$ the contracted vertex corresponding to $u$ and $x$ in $G/f$. First observe that as $e$ and $f$ share an endpoint
${\rm R}_{\rm eff}^{G_{n,p}/f}(\tilde{u}\leftrightarrow v)={\rm R}_{\rm eff}^{G_{n,p}}(v\leftrightarrow\{u,x\}).$
Applying Lemma 3.3 for the vertices $u,v,x$ with (3.8) then gives that for some fixed $C^{\prime}>0$
$\Big{|}{\rm R}_{\rm eff}(v\leftrightarrow\{u,x\})-\frac{3}{2pn}\Big{|}\leq C^{\prime}\frac{\epsilon}{pn},$
completing the proof by possibly enlarging the constant $C$. ∎
The bounds on the effective resistance for $G_{n,p}$ almost give the correct bounds on the effective resistance of $(G,\textup{{w}})$ by letting $p=1/\beta\gg\log n/n$ and lowering the weights of the edges in the coupling to $\exp(-\beta p)$, so that
${\rm R}_{\rm eff}(u\leftrightarrow v)\leq\frac{(2+o(1))}{np}e^{\beta p}\approx\frac{2\cdot e}{n\mu}.$
We shall improve the constant using the following straightforward lemma.
Lemma 3.4.
Let $\vartheta_{1},\ldots,\vartheta_{m}$ be unit flows (cf. Section 2.1) from $u$ to $v$ that are supported on disjoint edges, that is for each edge $e$ we have $\vartheta_{i}(e)>0$ for at most one $i\in\{1,\ldots,m\}$. Then, for any $\alpha_{1},\ldots,\alpha_{m}\geq 0$ we have
$\Big{\|}\sum_{i=1}^{m}\alpha_{i}\vartheta_{i}\Big{\|}^{2}_{r}=\sum_{i=1}^{m}\alpha^{2}_{i}\|\vartheta_{i}\|^{2}_{r}.$
Proof.
As the support of the flows are disjoint, we have that
$\Big{\|}\sum_{i=1}^{m}\alpha_{i}\vartheta_{i}\Big{\|}^{2}_{r}=\frac{1}{2}\sum_{e\in E}\Big{(}\sum_{i=1}^{m}\alpha_{i}\vartheta_{i}(e)\Big{)}^{2}=\frac{1}{2}\sum_{e\in E}\sum_{i=1}^{m}\alpha_{i}^{2}\vartheta_{i}(e)^{2}=\sum_{i=1}^{m}\alpha^{2}_{i}\|\vartheta_{i}\|^{2}_{r}.\qed$
We are now ready to prove the following lemma.
Lemma 3.5.
Assume $\beta=\beta(n)\ll n/\log n$ and fix $K>0$. Then the following holds with probability at least $1-O(n^{-K})$:
  1. (i)
    for all $u,v\in V$
    ${\rm R}_{\rm eff}(u\leftrightarrow v)\leq\frac{2(1+o(1))}{n\mu};$ (3.9)
  2. (ii)
    and for all edges $e=(u,v)$, $f=(u,x)$ with $|e\cap f|=1$
    ${\rm R}_{\rm eff}^{G/f}(u\leftrightarrow v)\leq\frac{3(1+o(1))}{2n\mu}.$ (3.10)
Proof.
Consider a coupling to a random graph similar to the one in Section 2.2 by letting $G_{n}[p_{0},p_{1})$ be the subgraph consisting of edges satisfying
$e\in G_{n}[p_{0},p_{1})\iff\omega_{e}\in[p_{0},p_{1})$
for $0\leq p_{0}\leq p_{1}\leq 1$. Define $p=\kappa(n)\log n/n$ for some function $\kappa=\kappa(n)\rightarrow\infty$ such that $\beta=o(p^{-1})$. Let $m:=\lfloor p^{-1}\rfloor$, and consider the sequence of snapshots
$p_{j}=j\cdot p\qquad j=1,\ldots,m.$
We will construct a flow on $G$ by considering a weighted sum of disjoint flows on $G_{n}[p_{j-1},p_{j})$. Namely, for each $j$ let $\vartheta_{j}$ be the current unit flow (the unique flow minimizing the energy in Definition 2.1) on $G_{n}[p_{j-1},p_{j})$, and let
$\tilde{\alpha}_{j}=e^{-\beta\cdot p_{j}}=e^{-j\beta p},\qquad j$ $=1,\ldots,m,$
$\alpha_{j}=\frac{\tilde{\alpha}_{j}}{\sum_{i=1}^{m}\tilde{\alpha}_{i}}=\tilde{\alpha}_{j}\frac{e^{\beta p}-1}{1-e^{-\beta mp}},\qquad j$ $=1,\ldots,m.$
As $\sum_{i=j}^{m}\alpha_{j}=1$, the flow defined by
$\vartheta:=\sum_{j=1}^{m}\alpha_{j}\vartheta_{j}$
is a unit flow from $u$ to $v$ where each $\vartheta_{j}$ is supported on different edges. Lemma 3.4 then gives that
${\rm R}_{\rm eff}(u\leftrightarrow v)\leq\|\vartheta\|^{2}_{r}=\sum_{j=1}^{m}\alpha^{2}_{j}\|\vartheta_{j}\|^{2}_{r}.$ (3.11)
By Rayleigh’s monotonicity principle, the effective resistance on $G_{n}[p_{j-1},p_{j})$ is upper bounded by the effective resistance on the same graph where each weight is decreased to $\exp(-j\beta p)$. Thus by inequality (3.6) of Lemma 3.2 (with say $\epsilon=\kappa^{-1/4}$) we have
$\|\vartheta_{j}\|^{2}_{r}\leq\frac{2(1+o(1))}{pn}e^{j\beta p}.$
for all $j$ with probability at least $1-O(n^{-K})$. Inserting this bound into (3.11) for each $j$ gives that
${\rm R}_{\rm eff}(u\leftrightarrow v)$ $\leq\frac{2(1+o(1))}{pn}\sum_{j=1}^{m}\frac{\tilde{\alpha}_{j}^{2}}{\big{(}\sum_{j=1}^{m}\tilde{\alpha}_{j}\big{)}^{2}}e^{j\beta p}$
$=\frac{2(1+o(1))}{pn}\Big{(}\frac{e^{\beta p}-1}{1-e^{-\beta mp}}\Big{)}.$
As $mp\geq 1-p$ and $\beta p=o(1)$, a Taylor expansion in the last line finishes the proof of inequality (3.9). The proof of (3.10) follows along the same lines by using inequality (3.7) in Lemma 3.2 instead of the inequality in (3.6). ∎
We may now easily prove Theorem 1.1 for the low disorder regime of $\beta\ll n/\log n$.
Proof of (1.3) in Theorem 1.1.
By equation (2.1), Lemma 3.1 and Lemma 3.5 we have that for $\beta\ll n/\log n$
$\mathcal{O}(\beta)=(1+o(1))\frac{4}{n^{2}\mu^{2}}\sum_{e\in E}\textup{{w}}(e)^{2}$
with high probability. Using the concentration bound of (3.1) for the sum (with $\beta^{\prime}=2\beta$, $m=|E|=n(n-1)/2$ and, say, $\delta=n^{-1/4}$) finishes the proof. ∎
3.3. Transfer-impedance
Recall the definition of the transfer-impedance matrix from Section 2.1 and that $|e\cap f|$ was defined to be the number of endpoints that edges $e$ and $f$ share. Let $Y^{0}$ be the transfer-impedance matrix in the unweighted case, that is
$|Y^{0}(e,f)|=\begin{cases}\frac{2}{n},&\text{if }|e\cap f|=2,\\ \frac{1}{n},&\text{if }|e\cap f|=1,\\ 0,&\text{if }|e\cap f|=0.\end{cases}$
In the following lemma, we give sharp bounds on the entries of the transfer-impedance matrix in the weighted case.
Lemma 3.6.
If $\beta=\beta(n)\ll n/\log n$, then for any $K>0$ and all edges $e,f$
$|Y(e,f)|=\frac{\textup{{w}}(f)}{n\mu}\big{(}|e\cap f|+o(1)\big{)}.$ (3.12)
with probability at least $1-O(n^{-K})$. In particular,
$\Big{|}Y(e,f)-\frac{\textup{{w}}(f)}{\mu}Y^{0}(e,f)\Big{|}=o(1)\cdot\frac{\textup{{w}}(f)}{n\mu}.$ (3.13)
with probability at least $1-O(n^{-K})$.
Proof.
Recall from Section 2.1 that
$Y(e,e)=\boldsymbol{\mathrm{P}}(e\in{\mathcal{T}})=\textup{{w}}(e){\rm R}_{\rm eff}(e^{-}\leftrightarrow e^{+}),$
and for $e,f$ with $|e\cap f|\leq 1$
$Y(e,f)^{2}$ $=\frac{\textup{{w}}(f)}{\textup{{w}}(e)}\big{(}\boldsymbol{\mathrm{P}}(e\in{\mathcal{T}})\boldsymbol{\mathrm{P}}(f\in{\mathcal{T}})-\boldsymbol{\mathrm{P}}(e,f\in{\mathcal{T}})\big{)}$
$=\textup{{w}}(f)^{2}{\rm R}_{\rm eff}(f^{-}\leftrightarrow f^{+})\Big{(}{\rm R}_{\rm eff}(e^{-}\leftrightarrow e^{+})-{\rm R}_{\rm eff}^{G/f}(e^{-}\leftrightarrow e^{+})\Big{)}$ (3.14)
Rayleigh’s monotonicity principle shows that for edges with $|e\cap f|=0$ we have
${\rm R}_{\rm eff}^{G/f}(e^{-}\leftrightarrow e^{+})\leq{\rm R}_{\rm eff}(e^{-}\leftrightarrow e^{+}),$
so that the bounds bounds from Lemma 3.1 and Lemma 3.5 give for $|e\cap f|\leq 1$
${\rm R}_{\rm eff}(e^{-}\leftrightarrow e^{+})$ $=(1+o(1))\frac{2}{n\mu},$
${\rm R}_{\rm eff}^{G/f}(e^{-}\leftrightarrow e^{+})$ $=(1+o(1))\Big{(}\frac{4-|e\cap f|}{2n\mu}\Big{)},$
with probability at least $1-O(n^{-K})$. Inserting this back into (3.14) readily gives (3.12).
For equation (3.13), notice that the signs of $Y(e,f)$ and $Y^{0}(e,f)$ agree as long as $|e\cap f|\geq 1$. Indeed, any edge pointing away from $e^{-}$ or towards $e^{+}$ can only have non-negative flow, while flow is non-positive for edges pointing towards $e^{-}$ or away from $e^{+}$. If $|e\cap f|=0$, then $Y^{0}(e,f)=0$, and hence the equality of (3.12) implies the equality of (3.13) regardless of $|e\cap f|$. ∎
Denote by $\boldsymbol{\mathrm{P}}^{0}$ the unweighted UST measure on $G$. Lemma 3.6 has the following corollary.
Corollary 3.7.
Fix $k\geq 1$ and let $e_{1},\ldots,e_{k}\in E$ be distinct edges. If $\beta=\beta(n)\ll n/\log n$, then for any $K>0$
$\boldsymbol{\mathrm{P}}(e_{1},\ldots,e_{k}\in{\mathcal{T}})=(1+o(1))\prod_{i=1}^{k}\frac{\textup{{w}}(e_{i})}{\mu}\boldsymbol{\mathrm{P}}^{0}(e_{1},\ldots,e_{k}\in{\mathcal{T}}).$
with probability at least $1-O(n^{-K})$.
Proof.
W.l.o.g. assume that $e_{1},\ldots,e_{k}$ contains no cycles. By the transfer-impedance theorem (Theorem 2.4) and Lemma 3.6, we have with probability $1-O(n^{-K})$ that
$\boldsymbol{\mathrm{P}}(e_{1},\ldots,e_{k}\in{\mathcal{T}})$ $=\sum_{\sigma\in S_{k}}\Big{(}sgn(\sigma)\prod_{i=1}^{k}Y(e_{i},e_{\sigma(i)})\Big{)}$
$=\sum_{\sigma\in S_{k}}\bigg{[}sgn(\sigma)\prod_{i=1}^{k}\frac{\textup{{w}}(e_{i})}{\mu}\prod_{i=1}^{k}\Big{(}Y^{0}(e_{i},e_{\sigma(i)})+\frac{o(1)}{n}\Big{)}\bigg{]}$
$=\prod_{i=1}^{k}\frac{\textup{{w}}(e_{i})}{\mu}\bigg{[}\Big{(}\sum_{\sigma\in S_{k}}sgn(\sigma)\prod_{i=1}^{k}Y^{0}(e_{i},e_{\sigma(i)})\Big{)}+k!\frac{k\cdot o(1)}{n^{k}}\bigg{]}$
$=(1+o(1))\prod_{i=1}^{k}\frac{\textup{{w}}(e_{i})}{\mu}\boldsymbol{\mathrm{P}}^{0}(e_{1},\ldots,e_{k}\in{\mathcal{T}}),$
where we used that for fixed $k$
$\boldsymbol{\mathrm{P}}^{0}(e_{1},\ldots,e_{k}\in{\mathcal{T}})\geq\frac{1}{k!n^{k}}\gg k!\frac{k\cdot o(1)}{n^{k}},$ (3.15)
so that we can absorb the $o(1)\cdot k!kn^{-k}$ error term into $o(1)$ times the probability of the edges being in the unweighted UST. ∎
We remark that in view of (3.15) to make Corollary 3.7 useful we must pick $K>k$.
3.4. Proof of Theorem 1.3 for small $\beta$ via tree moments
For random variables $X_{n}$ taking values in $\mathbb{R}$, the method of moments says that if all moments of $X_{n}$ converge to the moments of $X$ and these moments do not grow too fast, then $X_{n}$ converges in distribution to $X$. In [4], the authors have generalized this result to $t$th tree moments and local convergence of trees. Namely, for a finite rooted tree $(t,s_{1})$ and a rooted graph $(G,s_{2})$, a map $f$ from $V(t)$ to $V(G)$ is called a tree-map if $f$ is injective, maps $s_{1}$ to $s_{2}$, and if $u\sim v$ in $t$ then $f(u)\sim f(v)$ in $G$. Denote by $N(G,t)$ the number of distinct tree-maps between $(t,s_{1})$ and $(G,s_{2})$. Let $(\boldsymbol{\mathrm{P}}_{n})_{n\geq 1}$ and $\boldsymbol{\mathrm{P}}_{\infty}$ be probability distributions on rooted trees with corresponding expectations $(\boldsymbol{\mathrm{E}}_{n})_{n\geq 1}$ and $\boldsymbol{\mathrm{E}}_{\infty}$. The following is an analog of the method of moments for finite trees.
Proposition 3.8 (Proposition 8.7 in [4]).
[…]
One such tree that is uniquely determined by its $t$th moments is $\mathcal{P}$ the Poisson(1) tree conditioned to survive forever (see Section 8 of [4]). Consequently, to prove the local convergence of ${\mathcal{T}}$ in the case $\beta\ll n/\log n$, it suffices to prove convergence of the expectations of $N({\mathcal{T}},t)$.
Proof of (1.5) in Theorem 1.3.
As in the proof of Theorem 5.3 in [4], for any finite rooted tree $t$ let $\phi(f)$ be the image (in $G$ rooted at $1$) of the tree map $f$. Let $K>|t|$, then by Corollary 3.7
$\mathbb{E}\big{[}\boldsymbol{\mathrm{P}}(\phi(f)\subseteq{\mathcal{T}})\big{]}=\big{(}1+o(1)\big{)}\boldsymbol{\mathrm{P}}^{0}\big{(}\phi(f)\subseteq{\mathcal{T}}\big{)}+O(n^{-K})=\big{(}1+o(1)\big{)}\boldsymbol{\mathrm{P}}^{0}\big{(}\phi(f)\subseteq{\mathcal{T}}\big{)}.$
Hence, summing over all tree maps $f$
$\mathbb{E}\Big{[}\boldsymbol{\mathrm{E}}[N({\mathcal{T}},t)]\Big{]}=\mathbb{E}\Big{[}\sum_{f}\boldsymbol{\mathrm{P}}\big{(}\phi(f)\subseteq{\mathcal{T}}\big{)}\Big{]}=\big{(}1+o(1)\big{)}\sum_{f}\boldsymbol{\mathrm{P}}^{0}\big{(}\phi(f)\subseteq{\mathcal{T}}\big{)},$
which converges to $\boldsymbol{\mathrm{E}}[N(\mathcal{P},t)]$ by Theorem 5.3 of [4]. Applying Proposition 3.8 finishes the proof. ∎
4. High disorder
In this section, we will first prove the edge overlap result in Theorem 1.1 for $\beta\gg n(\log n)^{2}$. Then we will adapt these techniques to show the convergence of the local limit. The key argument is that, with high probability, edges in the MST with weights outside a window of width $2\epsilon/n$ (for some $\epsilon=o(1)$ depending on $\beta$) must also appear in the RSTRE. We remark that here high disorder refers to $\beta\gg n$ and not to the high disorder regime $\beta\gg n^{4/3}$ as in [10].
4.1. Overlap
The edge overlap result in the high disorder regime is restated in the next lemma.
Lemma 4.1.
Let $\beta=\beta(n)\gg n(\log n)^{2}$, then with high probability
$\mathcal{O}(\beta)=(1-o(1))n.$ (4.1)
Proof.
Suppose $e=(u,v)\in E$ such that $e$ is an edge in the MST. Our proof will construct several different (high probability) events on which the probability of $e$ also being in the RSTRE is close to $1$. Define $p=\omega_{e}$ and $p^{-}:=\omega_{e}-e^{-n}$, so that with high probability the only edge added in the graph process $G_{n,p}$ (coupled to $\omega$ as in Section 2.2) between $p^{-}$ and $p$ is the edge $e$, i.e. the event
$\mathcal{G}_{e}:=\big{\{}G_{n,p}=G_{n,p^{-}}\cup\{e\}\big{\}}$ (4.2)
occurs with probability, say, at least $1-O(n^{-4})$. Let $\epsilon\geq n^{-1/4}$ be a sequence tending to zero (e.g. $\epsilon=1/\log\log n$), and consider the following two events
$\mathcal{M}_{e}$ :=$:=\big{\{}e\in\textrm{MST}\big{\}}\cap\big{\{}\omega_{e}\leq\frac{5\log n}{n}\big{\}},$ (4.3)
$\mathcal{S}_{e}=\mathcal{S}_{e}(\epsilon)$ :=$:=\big{\{}|\mathcal{C}_{2}(p^{-})|\leq L\epsilon^{-2}\log n\big{\}},$ (4.4)
for some large enough but fixed constant $L$ as in Theorem 2.5. By Theorem 2.6 the graph $G_{n,5\log n/n}$ is connected with very high probability so that
$\mathbb{P}(\mathcal{M}_{e})=\frac{2}{n}-O(n^{-4}).$
Let
$E^{*}:=\Big{\{}e\in E\,:\,\omega_{e}\not\in\Big{[}\frac{1-\epsilon}{n},\frac{1+\epsilon}{n}\Big{]}\Big{\}},$
then using bounds on the size of the 2nd largest cluster of Theorem 2.5 gives that
$\mathbb{P}(\mathcal{S}_{e})\geq\mathbb{P}\big{(}\mathcal{S}_{e}\mid e\in E^{*}\big{)}\big{(}1-\frac{2\epsilon}{n}\big{)}=1-\frac{2\epsilon}{n}-O(n^{-\eta}).$ (4.5)
For the moment restrict to the events $\mathcal{M}_{e}$, $\mathcal{S}_{e}$ and $\mathcal{G}_{e}$. Let $S(u)$ and $S(v)$ be the connected components of the random graph process at $p^{-}$, and w.l.o.g. assume that $|S(u)|\leq|S(v)|$. Applying Kirchhoff’s formula together with the Nash-Williams inequality (see [8, Section 2.5]) using the cutset $E(S(u),S(u)^{c})$ gives that
$\boldsymbol{\mathrm{P}}(e\in{\mathcal{T}})\geq\frac{\textup{{w}}(e)}{\sum_{f\in E(S(u),S(u)^{c})}\textup{{w}}(f)}=\frac{1}{1+\sum_{f\in E(S(u),S(u)^{c})\setminus\{e\}}e^{-\beta(\omega_{f}-\omega_{e})}}.$ (4.6)
Notice that on $\mathcal{M}_{e}$ and $\mathcal{G}_{e}$ the edge $e$ has the lowest weight among all edges in $E(S(u),S(u)^{c})$. Furthermore, the random variables in the collection
$\Big{(}\frac{\omega_{f}-\omega_{e}}{1-\omega_{e}}\Big{)}_{f\in E(S(u),S(u)^{c})\setminus\{e\}}$
are i.i.d. with law uniform on $[0,1]$, and, on the event $\mathcal{S}_{e}$, the cutset is bounded in size by
$\big{|}E(S(u),S(u)^{c})\big{|}\leq n|\mathcal{C}_{2}(p)|\leq Ln\epsilon^{-2}\log n.$
Further, define
$\mathcal{B}_{e}$ :=$:=\Big{\{}\forall f\in E\big{(}S(u),S(u)^{c}\big{)}\setminus\{e\}\,:\,e^{-\beta(\omega_{f}-\omega_{e})}<\frac{1}{n^{5}}\Big{\}}$
=$=\Big{\{}\forall f\in E\big{(}S(u),S(u)^{c}\big{)}\setminus\{e\}\,:\,\frac{\omega_{f}-\omega_{e}}{1-\omega_{e}}>5\frac{\log n}{\beta(1-\omega_{e})}\Big{\}}$ (4.7)
As $\omega_{e}\leq 5\log n/n$, it follows that for $U$ distributed uniformly on $[0,1]$ that, on $\mathcal{M}_{e}$, $\mathcal{S}_{e}$ and $\mathcal{G}_{e}$, we have
$\mathbb{P}\big{(}\mathcal{B}_{e}\big{)}\geq\mathbb{P}\Big{(}U>6\frac{\log n}{\beta}\Big{)}^{Ln\epsilon^{-2}\log n}=\Big{(}1-6\frac{\log n}{\beta}\Big{)}^{Ln\epsilon^{-2}\log n}.$
Hence, if for some $\gamma>0$
$\frac{n\epsilon^{-2}(\log n)^{2}}{\beta}\leq\gamma,$ (4.8)
then the event $\mathcal{B}_{e}$ occurs with probability at least $1-O(\gamma)$.
Notice that inequality (4.6) gives that restricted to the events $\mathcal{M}_{e}$, $\mathcal{S}_{e}$, $\mathcal{G}_{e}$ and $\mathcal{B}_{e}$, we have
$\boldsymbol{\mathrm{P}}(e\in{\mathcal{T}})\geq\frac{1}{1+n^{-3}}\geq 1-n^{-3}.$ (4.9)
Hence, combing the previous bounds shows that for $\gamma\geq n^{-\eta}$ (where $\eta$ is the constant from Theorem 2.5)
$\mathbb{E}\big{[}\boldsymbol{\mathrm{P}}(e\in{\mathcal{T}})^{2}\big{]}\geq(1-n^{-3})^{2}\mathbb{P}\big{(}\mathcal{B}_{e}\mid\mathcal{M}_{e},\mathcal{S}_{e},\mathcal{G}_{e}\big{)}\mathbb{P}\big{(}\mathcal{M}_{e},\mathcal{S}_{e},\mathcal{G}_{e}\big{)}\geq\big{(}1-O(\gamma)\big{)}\big{(}1-\epsilon\big{)}\frac{2}{n},$
and in particular
$\mathbb{E}\big{[}\mathcal{O}(\beta)\big{]}\geq\big{(}1-O(\gamma)\big{)}\big{(}1-\epsilon\big{)}n.$
If $\beta\gg n(\log n)^{2}$, then we may pick $\gamma=o(1)$ and $\epsilon=o(1)$ such that the condition (4.8) is fulfilled. Furthermore, as trivially the second moment of $\mathcal{O}(\beta)$ is bounded from above by $n^{2}$, applying Chebyshev’s inequality immediately gives (4.1). ∎
4.2. Local limit
To show convergence of the local limit, we will construct similar events as in the proof of Lemma 4.1 for all outgoing edges of a fixed vertex. We first consider the case $r=1$ in the following lemma, and then extend this to the full local limit by using the union bound and transitivity. Recall that $B_{\mathcal{T}}(u,r)$ and $B_{M}(u,r)$ are the closed balls of radius $r$ centered at $u$ in the RSTRE and MST respectively.
Lemma 4.2.
For any $u\in V$ we have that
$\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(u,1)\neq B_{M}(u,1)\big{)}\Big{]}=O\Bigg{(}\max\Bigg{\{}\Big{(}\frac{n(\log n)^{5}}{\beta}\Big{)}^{1/3},n^{-\eta/2}\Bigg{\}}\Bigg{)},$ (4.10)
where $\eta$ is the constant from Theorem 2.5.
Proof.
We first show that the expectation of $\boldsymbol{\mathrm{P}}(B_{M}(u,1)\not\subseteq B_{\mathcal{T}}(u,1))$ is small by restricting to the events as in the proof of Lemma 4.1. Let
$\gamma=\epsilon=\max\Big{\{}\Big{(}\frac{n(\log n)^{2}}{\beta}\Big{)}^{1/3},n^{-2\eta/3}\Big{\}}.$
Our choice of $\epsilon$ and $\gamma$ guarantees that the condition in (4.8) holds (where we may implicitly assume that $\epsilon=o(1)$). For $e=(u,v)$ recall the definitions of $\mathcal{G}_{(u,v)}$, $\mathcal{S}_{(u,v)}(\epsilon)$ and $\mathcal{B}_{(u,v)}$ from (4.2), (4.4) and (4.7), and further define
$\mathcal{C}$ :=$:=\big{\{}G_{n,5\log n/n}\textrm{ is connected}\big{\}},$
$\mathcal{I}$ $:=\Big{[}0,\frac{1-\epsilon}{n}\Big{]}\cup\Big{[}\frac{1+\epsilon}{n},5\frac{\log n}{n}\Big{]},$
$\mathcal{E}$ :=$:=\big{\{}\forall(u,v)\in M:\omega_{(u,v)}\in\mathcal{I}\big{\}}.$
By a union bound and we have that
$\boldsymbol{\mathrm{P}}\big{(}B_{M}(u,1)\not\subseteq B_{\mathcal{T}}(u,1)\big{)}\leq\sum_{v\neq u}\boldsymbol{\mathrm{P}}\big{(}(u,v)\not\in{\mathcal{T}}\big{)}1_{(u,v)\in M},$
and by the arguments leading to (4.9)
$\sum_{v\neq u}\boldsymbol{\mathrm{P}}\big{(}(u,v)\not\in{\mathcal{T}}\big{)}1_{(u,v)\in M}1_{\mathcal{G}_{(u,v)}}1_{\mathcal{S}_{(u,v)}}1_{\mathcal{B}_{(u,v)}}1_{\mathcal{C}}\leq n^{-2}.$ (4.11)
We now split the expectation into different terms depending on whether the events of the indicators in (4.11) hold or not. First, restrict to the events $\mathcal{G}_{(u,v)}$, $\mathcal{C}$ and $\mathcal{E}$ such that with the bound around (4.2) and Theorem 2.6, we have
$\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{M}(u,1)\not\subseteq B_{\mathcal{T}}(u,1)\big{)}\Big{]}$ $\leq\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{M}(u,1)\not\subseteq B_{\mathcal{T}}(u,1)\big{)}1_{\mathcal{C}}1_{\mathcal{E}}\Big{]}+\mathbb{P}(\mathcal{E}^{c}\cap\mathcal{C})+O(n^{-4})$
$\leq\mathbb{E}\Big{[}\sum_{v\neq u}\boldsymbol{\mathrm{P}}\big{(}(u,v)\not\in{\mathcal{T}}\big{)}1_{(u,v)\in M}1_{\mathcal{G}_{(u,v)}}1_{\mathcal{C}}1_{\mathcal{E}}\Big{]}+O(\epsilon),$
where, using the coupling of the MST to the random graph, we may bound
$\mathbb{P}(\mathcal{E}^{c}\cap\mathcal{C})\leq\mathbb{P}\Big{(}\exists v\neq u:\omega_{(u,v)}\in\Big{[}\frac{1-\epsilon}{n},\frac{1+\epsilon}{n}\Big{]}\Big{)}=O(\epsilon).$
Restricting further to $\mathcal{S}_{(u,v)}$ we obtain
$\mathbb{E}\Big{[}\sum_{v\neq u}\boldsymbol{\mathrm{P}}\big{(}(u,v)\not\in{\mathcal{T}}\big{)}1_{(u,v)\in M}1_{\mathcal{G}_{(u,v)}}1_{\mathcal{C}}1_{\mathcal{E}}\Big{]}$ $\leq\mathbb{E}\Big{[}\sum_{v\neq u}\boldsymbol{\mathrm{P}}\big{(}(u,v)\not\in{\mathcal{T}}\big{)}1_{(u,v)\in M}1_{\mathcal{G}_{(u,v)}}1_{\mathcal{S}_{(u,v)}}1_{\mathcal{C}}\Big{]}$
$\qquad+\sum_{v\neq u}\mathbb{P}\Big{(}(u,v)\in M,\mathcal{S}_{(u,v)}^{c},\mathcal{E}\Big{)}.$
The bound from (4.5) (see also Theorem 2.5) gives us that
$\sum_{v\neq u}\mathbb{P}\Big{(}(u,v)\in M,\mathcal{S}_{(u,v)}^{c},\mathcal{E}\Big{)}\leq n\mathbb{P}\big{(}\mathcal{S}_{(u,v)}^{c}\mid\omega_{(u,v)}\in I\big{)}\mathbb{P}\big{(}\omega_{(u,v)}\in I\big{)}=O(n^{-\eta}\log n).$
Finally, restricting to $\mathcal{B}_{(u,v)}$ gives us with the condition (4.8) that
$\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{M}(u,1)\not\subseteq B_{\mathcal{T}}(u,1)\big{)}\Big{]}$ $\leq\mathbb{E}\Big{[}\sum_{v\neq u}\boldsymbol{\mathrm{P}}\big{(}(u,v)\not\in{\mathcal{T}}\big{)}1_{(u,v)\in M}1_{\mathcal{G}_{(u,v)}}1_{\mathcal{S}_{(u,v)}}1_{\mathcal{B}_{(u,v)}}1_{\mathcal{C}}\Big{]}$
$+\sum_{v\neq u}\mathbb{P}\Big{(}\mathcal{B}_{e}^{c}\mid(u,v)\in M,\mathcal{S}_{(u,v)},\mathcal{G}_{(u,v)},\mathcal{C}\Big{)}\mathbb{P}\big{(}(u,v)\in M\big{)}+O(\epsilon)$
$\leq n^{-2}+nO(\gamma)\cdot\frac{2}{n}+O(\epsilon)=O(\gamma).$ (4.12)
Consider now the degrees $\deg_{M}(u)=|B_{M}(u,1)|$ and $\deg_{\mathcal{T}}(u)=|B_{\mathcal{T}}(u,1)|$ of $u$ in $M$ and ${\mathcal{T}}$, respectively. Both degrees have expectation $2(n-1)/n$, and by the calculations above, the degree $\deg_{\mathcal{T}}(u)$ almost stochastically dominates $\deg_{M}(u)$. We will use this fact now to show that the ball in ${\mathcal{T}}$ is with high probability not larger than the ball in $M$.
By (4.12) we have
$\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(u,1)\not\subseteq B_{M}(u,1)\big{)}\Big{]}$ $\leq\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(u,1)\not\subseteq B_{M}(u,1),B_{M}(u,1)\subseteq B_{\mathcal{T}}(u,1)\big{)}\Big{]}$
$\qquad+\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{M}(u,1)\not\subseteq B_{\mathcal{T}}(u,1)\big{)}\Big{]}$
$\leq\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}\deg_{\mathcal{T}}(u)>\deg_{M}(u)\big{)}\Big{]}+O(\gamma).$ (4.13)
Notice that (4.12) immediately also shows that
$\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}\deg_{\mathcal{T}}(u)<\deg_{M}(u)\big{)}\Big{]}=O(\gamma).$
Therefore, considering the cases where $\deg_{\mathcal{T}}(u)$ is larger than $\deg_{M}(u)$ or not, gives
00 0 $=\mathbb{E}\Big{[}\boldsymbol{\mathrm{E}}\big{[}\deg_{\mathcal{T}}(u)-\deg_{M}(u)\big{]}\Big{]}$
$=\mathbb{E}\Big{[}\boldsymbol{\mathrm{E}}\big{[}\big{(}\deg_{\mathcal{T}}(u)-\deg_{M}(u)\big{)}1_{\deg_{\mathcal{T}}(u)>\deg_{M}(u)}\big{]}\Big{]}$
$\qquad\qquad+\mathbb{E}\Big{[}\boldsymbol{\mathrm{E}}\big{[}\big{(}\deg_{\mathcal{T}}(u)-\deg_{M}(u)\big{)}1_{\deg_{\mathcal{T}}(u)<\deg_{M}(u)}\big{]}\Big{]}$
$\geq\mathbb{E}\big{[}\boldsymbol{\mathrm{P}}\big{(}\deg_{\mathcal{T}}(u)>\deg_{M}(u)\big{)}\big{]}-60\log n\mathbb{E}\big{[}\boldsymbol{\mathrm{P}}\big{(}\deg_{\mathcal{T}}(u)<\deg_{M}(u)\big{)}\big{]}$
$\qquad\qquad-n\mathbb{P}\big{(}\deg_{M}(u)>60\log n\big{)}.$ (4.14)
Rearranging (4.14) and using Theorem 2.6 shows that
$\mathbb{E}\big{[}\boldsymbol{\mathrm{P}}\big{(}\deg_{\mathcal{T}}(u)>\deg_{M}(u)\big{)}\big{]}=O(\gamma\log n).$ (4.15)
Putting (4.12), (4.13) and (4.15) together finishes the proof. ∎
As a corollary of Lemma 4.2 we may now obtain the local limit of the RSTRE for $\beta=n(\log n)^{\lambda(n)}$ for diverging $\lambda(n)$.
Proof of (1.6) in Theorem 1.3.
Fix $r\geq 1$. If $B_{M}(1,r)\neq B_{\mathcal{T}}(1,r)$, then there must exist at least one vertex $u$ in $B_{M}(1,r)$ such that $B_{M}(u,1)\neq B_{\mathcal{T}}(u,1)$. A union bound gives that
$\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(1,r)\neq B_{M}(1,r)\big{)}\Big{]}$ $\leq\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(1,1)\neq B_{M}(1,1)\big{)}\Big{]}$
$\qquad+\sum_{u\neq 1}\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(u,1)\neq B_{M}(u,1)\big{)}1_{u\in B_{M}(1,r)}\Big{]}.$ (4.16)
Using exchangeability of the edge weights (due to transitivity of the complete graph) and the fact that $u\in B_{M}(1,r)$ if and only if $1\in B_{M}(u,r)$, gives for any $v\neq u$ that
$\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(u,1)\neq B_{M}(u,1)\big{)}1_{u\in B_{M}(1,r)}\Big{]}$ $=\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(u,1)\neq B_{M}(u,1)\big{)}1_{1\in B_{M}(u,r)}\Big{]}$
$=\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(u,1)\neq B_{M}(u,1)\big{)}1_{v\in B_{M}(u,r)}\Big{]}.$
It follows that
$\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(u,1)\neq B_{M}(u,1)\big{)}1_{u\in B_{M}(1,r)}\Big{]}$
$\qquad\qquad=\frac{1}{n-1}\sum_{v\neq u}\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(u,1)\neq B_{M}(u,1)\big{)}1_{v\in B_{M}(u,r)}\Big{]}$
$\qquad\qquad=\frac{1}{n-1}\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(u,1)\neq B_{M}(u,1)\big{)}\big{(}\big{|}B_{M}(u,r)\big{|}-1\big{)}\Big{]}$ (4.17)
In view of Lemma 2.7, the size of $|B_{M}(u,r)|$ is bounded with high probability by $60^{r}(\log n)^{r}$, so that (4.17) can be bounded by
$\frac{60^{r}(\log n)^{r}}{n-1}\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(u,1)\neq B_{M}(u,1)\big{)}\Big{]}+O(n^{-4}).$ (4.18)
Plugging this bound into (4.16), and using that $\beta\gg n(\log n)^{5+3r}$ shows with Lemma 4.2 that
$\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(1,r)\neq B_{M}(1,r)\big{)}\Big{]}=o(1).$
Consequently, for any finite rooted tree $t$ of height $r$
^$\widehat{\mathbb{P}}(B_{\mathcal{T}}(1,r)\simeq t)$ $\leq\mathbb{P}(B_{M}(1,r)\simeq t)+\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(1,r)\neq B_{M}(1,r)\big{)}\Big{]}$
$=\mathbb{P}(B_{M}(1,r)\simeq t)+o(1)$
and
^$\widehat{\mathbb{P}}(B_{\mathcal{T}}(1,r)\simeq t)$ $\geq\mathbb{P}(B_{M}(1,r)\simeq t)-\mathbb{E}\Big{[}\boldsymbol{\mathrm{P}}\big{(}B_{\mathcal{T}}(1,r)\neq B_{M}(1,r)\big{)}\Big{]}$
$=\mathbb{P}(B_{M}(1,r)\simeq t)-o(1).$
The local convergence of the MST on the complete graph in Theorem 2.9 then also implies the local convergence of the RSTRE (to the same limiting object). ∎
Remark 4.3.
We expect the high disorder local limit result to also hold when $\lambda=\lambda(n)$ is some large fixed constant instead of diverging to infinity. In (4.18) we used that, with high probability, the neighborhood $B_{M}(1,r)$ grows in size at most as fast as $60^{r}(\log n)^{r}$, however, as shown in [1], the typical size of a neighborhood is of order $r^{3}$. We are not aware of any better tail bounds, say $|B_{M}(1,r)|=O(\log n)$ for any fixed $r\geq 1$, that hold with large enough probability to make (4.18) work.
Acknowledgements
The author is grateful to Rongfeng Sun for helpful comments on a draft of this paper. Furthermore, the author would like to thank Michele Salvi for raising interesting questions about the RSTRE.